OpenStudy (anonymous):

What is the sum of the measures of the interior angles of a regular polygon where a single exterior angle measures 30°?

OpenStudy (experimentx):

solve for n (n-2)*180 = 180 - 30

OpenStudy (anonymous):

All regular polygons exterior angles add up to 30 degrees. 360/30 = 12 sides. Use formula (n-2) × 180 to find sum of interior angles. (12-2)x180 = 1800 total interior degrees or 150 degrees each angle.

OpenStudy (cwrw238):

divide 360 by 30 gives number of sides of polygon = 12 sides each interior angle = 180-30 = 150 total of int. angles = 12* 150

OpenStudy (experimentx):

sorry: (n-2)*180/n=150

OpenStudy (experimentx):

no ... that's the measure of each individual internal ange

OpenStudy (cwrw238):

no the sum of interior angles = 12 * 150 as there are 12 angles in the polygon and each int. angle = 150 degrees 12*150 = 1800 degres

OpenStudy (anonymous):

so it would be 1800

OpenStudy (cwrw238):

yes